OpenStudy (anonymous):

Please help and go through ll steps/show work...pretty please? Joanna invests her $5,000 bonus into a high yield account that earns 7.6% simple interest. She wants to buy a Burmese Rabbit Hound with the money. The breeder told her that a puppy would cost $5,500. How long will Joanna have to wait before she has $5,500 in her savings account?

OpenStudy (tkhunny):

Simple Interest: I = Prt Cost - Savings = Interest 5500 - 5000 = 500 500 = 5000*0.076*t Solove for t. The hard part is translating sentences to equations or expressions.

OpenStudy (anonymous):

so 380t=500 divide 380 by both sides and t=1.32?

OpenStudy (anonymous):

how many years would that translate to?

OpenStudy (tkhunny):

That is years. 500*0.076*1 = 380 -- One year! 500*0.076*2 = 760 -- Two years!

OpenStudy (anonymous):

thought i was suppose to divide?

OpenStudy (anonymous):

can you show the work and the answer? is it a fraction like 1 1/3 years or 16 months ?

OpenStudy (anonymous):

I'm sorry I am still confused

OpenStudy (tkhunny):

You did. I'm just trying to convince you that 't' is in years. If I were in a classroom, I would have walked over to teh other side of the room and my examples would not have been confusing.

OpenStudy (tkhunny):

500/380 = 1.3157894736842105263157894736842 1 Year 3 Months 3 Weeks 2 Days 0 Hours 37 Minutes 54 Seconds It's all the same. Why does anyone care what form you put it in?

OpenStudy (anonymous):

i know t is time or years but how does 1.32 convert to years?

OpenStudy (anonymous):

I'd say that it's all the same too but my instructor is very very picky she will find anything to knock points off

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I was just wondering because if there are 12 months in a year .3 is not 3 months its less?

OpenStudy (tkhunny):

Tough call. You just have to come to grips with t in years. If you get t = 1, that means one year. If you get t = 5, that means 5 years. If you get \(t = \sqrt{5}\), that means t = 2.23608 or so years. 1.31579 Years = 1 Year and Change Subtract the whole year. 0.31579 Years Multiply by 12, because there are 2 months in a year. 0.31579*12 = 3.78948 Months = 3 Months and Change Subtract the Whole Months 0.78948 Months Multiply by 25/6, because there is an average of \(4\dfrac{1}{6}\) weeks in the 12 months of the calendar. 0.78948 * (25/6) Weeks = 3.2895 Weeks = 3 Weeks and Change You just need the conversion factors.

OpenStudy (anonymous):

thanks for explaining it

OpenStudy (tkhunny):

Of course, I meant "because there are 12 months in a year"